Write the equation of a line parallel to the given line but passing through the given point.
y=-3x-6; (-1,5)

Answer by stanbon(75887)

Write the equation of a line parallel to the given line but passing through the given point.
y=-3x-6; (-1,5)
-----
Given equation slope = -3
-----
Parallel line slope must be -3
--------
Form: y = mx + b
5 = -3*-1+b
b = 5-3 = 2
-----
Equation:
y = -3x+2
=================
Cheers,
Stan H.
================